OpenStudy (cwtan):

How to integrate: \(\Huge \int x(2^x) dx\)?

OpenStudy (turingtest):

by parts\[dv=2^xdx\]\[u=x\]

OpenStudy (cwtan):

errrr..... how to integrate the 2^x? I know this can be done by part but i duno how to integrate 2^x...

OpenStudy (cruffo):

Are you still viewing this problem?

OpenStudy (anonymous):

The integral of 2^x is 2^x/ln2

OpenStudy (cwtan):

why and how......

OpenStudy (anonymous):

\[\int\limits\limits_{}^{}a^xdx=\int\limits\limits_{}^{}e^{\ln(a^x)}dx=\int\limits_{}^{}e^{xln(a)}dx=\frac{e^{xln(a)}}{\ln(a)}=\frac{a^x}{\ln(a)}\]

zepdrix (zepdrix):

You need to go back to logarithmic differentiate to understand that one :) \[y=a^x\]\[\ln y=x \ln a\]\[\frac{ y' }{ y }=\ln a\]\[y'=y \ln a = a^x \ln a\] So think about this a second, if you take the derivative of this function what do you get:\[y=\frac{ 1 }{ \ln a }a^x\]\[y'=(\frac{ 1 }{ \ln a })a^x \ln a = a^x\] Or writing it the way ivan did works really nicely too ^^ I like that better actually, heh.

OpenStudy (cwtan):

oh thanks for explaining!